### v1

# FAIR Enough: How Can We Develop and Assess a FAIR-Compliant Dataset for Large Language Model Training?

Abstract
Advancements in Large Language Models (LLMs) highlight the need for ethical practices and data integrity. We introduce a framework that embeds FAIR (Findable, Accessible, Interoperable, Reusable) data principles into LLM training. This approach marks a shift towards practices compliant with FAIR standards. Our framework presents guidelines for integrating FAIR data principles into LLM training. This initiative includes a checklist for researchers and developers. We also demonstrate its practical application through a case study focused on bias identification and mitigation in our FAIR-compliant dataset. This work is a significant contribution to AI ethics and data science, advocating for balanced and ethical training methods in LLMs.
Keywords Large Language Models, FAIR Principles, Data Management, AI Ethics, Interoperability, Dataset Curation, Bias
1 Introduction
In the dynamic realm of artificial intelligence (AI), Language Models (LMs) play a pivotal role in advancing diverse AI applications Jiang et al. (2020). From virtual assistants to content generation, LMs have become indispensable in both shaping the future trajectory of academic research and in their widespread practical applications. The impact of this transformation is further amplified by the advancements of Large Language Models (LLMs) Zhao et al. (2023b), such as OpenAI’s GPT series, Gemini, LLaMa, and Falcon. As of January 2024, LLM development has garnered \$18.2 billion in funding and \$2.1 billion in revenue, showcasing its substantial economic impact TrendFeedr (2024). The rapid success of LLMs underscores the importance of diverse data for broadening their applicability across different domains. This idea is also well-supported in research Bender et al. (2021); Wang et al. (2023), which affirms the importance of high-quality data in refining these models. The training of these big models on data from diverse sources highlights the complex ethical and responsible data practice challenges in their implementation Chang et al. (2023); Wang et al. (2023).
The FAIR data principles, which stand for Findable, Accessible, Interoperable, and Reusable Dunning et al. (1970); Boeckhout et al. (2018), were initially established to improve the stewardship of scientific data. These principles have gained increasing recognition in the field of responsible AI development Deshpande and Sharp (2022). In recent times, particularly in the context of generative AI, the importance of these principles has become more evident due to ethical challenges. These challenges include concerns about bias, privacy, and potential misuse of AI-generated content of Ethics (2024); Partescano et al. (2021). For LMs, including advanced LLMs  [^1] In this study, ’LLMs’ encompasses both ’LMs’ and ’LLMs’, with the latter representing advanced iterations of the former, adherence to FAIR data principles is even more crucial. The way data is curated and managed significantly affects the reliability, ethical integrity, and impact of these models.
The exploration of FAIR data principles has been pivotal in data science and management. Seminal works Dunning et al. (1970); Wilkinson et al. (2016); Hasnain and Rebholz-Schuhmann (2018); Jacobsen et al. (2020a); Shmueli et al. (2021) have provided extensive insights into these principles across various fields. However, their application in training and developing LLMs is an emerging and evolving area of research. Recent advances focus on aligning LLMs with ethical standards Chang et al. (2023); Wang et al. (2023), underscoring the importance of embedding FAIR principles in LLM training from the start. Specific challenges in LLMs, such as data biases Ntoutsi et al. (2020) and interpretability Singh et al. (2023), highlight the need for ethical and balanced training methods. This emphasis on data ethics is in line with broader trends in AI ethics and data science Raji et al. (2021); Jobin et al. (2019) stress the necessity of integrating these principles in LLM development.
We suggest a lifecycle model for developing LLMs that integrates FAIR principles at every phase. A key element of our methodology is developing a FAIR-compliant dataset that includes a range of narratives from various sources. Our main contribution extends beyond just data collection; it involves meticulously aligning this dataset with FAIR principles to improve LLM training effectiveness. While adhering to FAIR principles doesn’t guarantee optimal ethical compliance, it is a significant step in that direction. This research paves the way for further exploration in this area.The key contributions of our work are:
1.1 Contributions
  1. 1.
    We present a detailed review of FAIR data principles within AI research, including a comprehensive checklist for researchers and developers.
  2. 2.
    We introduce a novel framework that embeds FAIR data principles throughout the LLM training lifecycle, facilitating ethical and effective use in various AI contexts.
  3. 3.
    We demonstrate the practical benefits of a FAIR-compliant dataset through a case study. This study focuses on identifying and addressing biases [^2] Bias refers to a wide range, however, our case study particularly address linguistic biases, either explicit or implicit, targetting certain protected groups Fletcher et al. (2021); Garrido-Muñoz et al. (2021) before training LLMs.
1.2 Literature Selection Criteria
This study examines publications from the last five years, prioritized English-language articles from prominent AI and ML journals and conferences. We explored databases and proceedings from Elsevier, Springer, Nature portfolio, IEEE Transactions, and key conferences like NeurIPS and ICML. Recognizing the fast-paced evolution of LLM research, we included relevant preprints and seminal works for a comprehensive perspective.
Search Query:
("FAIR principles" OR "data management") AND "large language models"
AND "training data" AND "ethical considerations" FROM 2018 TO 2023 IN English
IN Journals and Conferences (e.g., JAIR, JMLR, IEEE Transactions, NeurIPS, ICLR).
This search resulted in 135 papers, out of which we carefully chose about 75 that directly pertained to FAIR data principles in the context of training models. We noted a scarcity of work combining LLMs with FAIR data principles, a gap that our study aims to address.
1.3 Comparative Analysis of Related Work
We have conducted a meta-review of key articles on FAIR data principles, categorizing them by domain and their relevance to LLMs. We also highlight the difference of our work compared to the previous works. The findings are summarized in Table 1.
Reference Summary Domain Focus LLM Relevance
Alvarez-Romero et al. (2023)
Discusses data governance for FAIR principles in health data management.
Health Data
Not Applicable
Inau et al. (2021)
Reviews initiatives for applying FAIR principles in managing health data.
Health Data
Not Applicable
Sadeh et al. (2023)
Examines FAIR data practices’ role in global mental health research.
Mental Health Research
Not Applicable
Stanciu (2023)
Analyzes FAIR principles in healthcare data, emphasizing cybersecurity.
Health Data
Partially (InstructGPT)
Raycheva et al. (2023)
Addresses challenges in FAIR application and legal aspects in European rare disease databases for ML technologies.
Health Data
Not Applicable
Vesteghem et al. (2019)
Elaborates on FAIR principles in precision oncology research.
Health Data
Not Applicable
Jacobsen et al. (2020b)
Offers a broad perspective on FAIR principles’ implementation across various fields.
Multi-Domain
Not Applicable
Our Study
Evaluates the application of FAIR data principles in the creation and training of datasets for LLMs. This study encompasses a range of models including the BERT and GPT families, as well as the LLaMa-2-7b model, highlighting diverse approaches in LLM development.
News Data
Yes (BERT/ , GPT families, LLaMa-2-7b)
Table 1: Summary of Review Papers on FAIR Data Principles
Table 1 highlights a predominant focus on health data in existing research. Notably, one study incorporates a LLM application (GPT). Our work, however, stands out for its application of FAIR data principles to a range of LLMs, including BERT-large, GPT, and LLaMa-2-7b. This approach broadens the scope and generalizability of our research.
2 FAIR Data Principles: Theoretical Background and Significance
The FAIR data principles significantly influence data management practices across various fields. These include physics Dungkek (2022), health Inau et al. (2023), environmental science Partescano et al. (2021), pharmaceuticals Wise et al. (2019), chemistry Jeliazkova et al. (2023), computer science Axton et al. (2016), research and development Wilcox (2018), and clinical studies Vesteghem et al. (2019). These principles ensure data is well-organized, accessible, and reusable in different research domains and applications. In AI and Natural Language Processing (NLP), the implementation of interdisciplinary strategies Huerta et al. (2023) and structured approaches Bernabé et al. (2023) is crucial to uphold ethical data practices Shmueli et al. (2021) and ensure fairness in AI models Bateni et al. (2022). Our detailed literature review elaborates on each FAIR principle, as depicted in Figure 1, and we also provide a comprehensive compliance checklist in Table 2333For hyperlinks to the technologies and APIs mentioned in the table, refer to Appendix .1.
Figure 1: FAIR Data Principles: Key Aspects of Findability, Accessibility, Interoperability, and Reusability in Data Management.
2.1 Findability
Findability, a crucial aspect of the FAIR data principles, ensures data and resources are easily locatable and accessible Findlay and Seah (2020). It involves creating detailed metadata, using persistent identifiers, and promoting efficient data indexing and searchability. This enhances discoverability by both humans and machines. Significant contributions to data findability within the FAIR framework include Bernabé et al. (2023), which explores strategies for FAIRification of data to enhance findability, and Santos et al. (2023), investigating a model for the FAIR digital object framework focused on digital objects’ findability. The application of FAIR principles in AI, particularly the importance of findability in AI datasets, is discussed in Huerta et al. (2023). Furthermore, Götz (2023) considers the trustworthiness of data repositories, emphasizing the critical role of findability in these environments.
2.2 Accessibility
Accessibility, as defined by the FAIR data principles, is the ease of obtaining and using data once located Boeckhout et al. (2018). It involves long-term data preservation, ethical access protocols, and ensuring data remains retrievable and usable over time. Discussions on trust in data repositories as part of accessibility are covered in Wang and Savard (2023), which also provides a comprehensive understanding of FAIR principles, including accessibility. The application of FAIR principles to research software, with a focus on accessibility, is explored in Lamprecht et al. (2020). How the Fedora platform supports FAIR data principles, particularly accessibility, is discussed in Wilcox (2018). Challenges and opportunities in implementing FAIR principles, including accessibility, in Brazilian data science are addressed in Sales et al. (2020).
2.3 Interoperability
Interoperability, as outlined in the FAIR data principles, refers to the ability of different data systems to work together efficiently da Silva Santos et al. (2016). It requires standardized data formats and protocols, enabling easy data exchange and integration across diverse systems. The Immune Epitope Database (IEDB) and its focus on interoperability in implementing FAIR principles are discussed in Vita et al. (2018). The eXtensible ontology development (XOD) principles and tools, which support interoperability in ontology development, are explored in He et al. (2018). Challenges and strategies for implementing FAIR data principles, including interoperability, in data science are addressed in Sales et al. (2020). A framework and metrics for evaluating the FAIRness of data, with a focus on interoperability, are provided in Wilkinson et al. (2018). The GO FAIR initiative, which aims to accelerate the adoption of FAIR principles including interoperability, is discussed in Schultes et al. (2018).
2.4 Reusability
Reusability, a key component of the FAIR data principles, ensures data is stored and documented for future retrieval and reuse Anguswamy and Frakes (2012). It encompasses rich metadata, legal and ethical considerations, and potential societal impact. A structured approach to planning data FAIRification, with an emphasis on reusability, is discussed in Bernabé et al. (2023). A model for digital objects under FAIR principles, focusing on reusability, is offered in Santos et al. (2023). The application of FAIR principles in AI, particularly regarding the reusability of AI datasets, is explored in Huerta et al. (2023). The importance of designing workflows with reusability as a primary goal is discussed in Wolf et al. (2021). A comprehensive exploration of FAIR principles, including detailed insights on reusability, is presented in Jacobsen et al. (2020a).
Table 2: Comprehensive Checklist for Ensuring FAIR Principles Compliance, Along with Associated Tools and Practices.
Principle Features Compliance list Tools and Practices
Findability
F1: Rich and descriptive metadata
F2: Standardized data indexing
F3: Documented data sources
F4: Advanced search functionalities
F5: Persistent identifiers like DOIs
F1: Metadata includes titles, authors, abstracts, keywords, and affiliations.
F2: Use of standardized taxonomy and ontology for indexing.
F3: Clear documentation of data sources and collection methodologies.
F4: Implementation of advanced search tools and interfaces.
F5: Assignment of persistent identifiers such as DOIs.
Metadata Management: Apache Atlas, Collibra, ORCID
Persistent Identifiers: CrossRef for DOIs
Data Indexing: Elasticsearch, Apache Solr, DSpace
Search Interfaces: Algolia, Apache Lucene
Data Repositories: NCBI, RE3data, CKAN, Dataverse, Zenodo, Figshare, EPrints, ResearchGate, Academia.edu
Accessibility
A1: Clear data access protocols
A2: Long-term data preservation
A3: Open access platforms
A4: Standardized APIs for data access
A5: Data availability in accessible formats
A1: Detailed access instructions and authentication processes.
A2: Use of reliable digital preservation services like CLOCKSS or Portico.
A3: Data deposited in open repositories such as Figshare or Zenodo.
A4: APIs conform to standards such as OpenAPI for ease of use.
A5: Data provided in multiple formats (e.g., CSV, JSON, XML) to ensure usability.
API Tools: OpenAPI, GraphQL, RESTful interfaces
Data Preservation: Archivematica, LOCKSS, Zenodo, Figshare
Cloud Storage: Amazon S3, Google Cloud Storage, Microsoft Azure
Ethical Access: OneTrust, TrustArc
Interoperability
I1: Standard data formats
I2: Common communication protocols
I3: Data exchange standards
I4: Tools for data transformation and mapping
I1: Data conforms to community-recognized standards (e.g., MIAME, Ecological Metadata Language).
I2: Support for protocols such as OAI-PMH for metadata harvesting.
I3: Use of frameworks like Schema.org for structured data.
I4: Availability of services like XSLT or OpenRefine for data conversion and mapping.
Data Formats: JSON, XML, CSV, DICOM, GenBank
Protocols: HTTP, SOAP, REST, gRPC
Data Standards: RDF, HL7 FHIR, ISO/IEC standards
Transformation Tools: XSLT, Talend, Informatica, Apache NiFi
Ontology Systems: OWL, SPARQL, XOD
Specialized Databases: IEDB
Reusability
R1: Detailed metadata for context
R2: Data curated for future use
R3: Adherence to ethical standards
R4: Licensing frameworks
R5: Consideration of societal impacts
R1: Comprehensive metadata including experimental conditions, methodologies, and provenance.
R2: Curating data with clear versioning and update records.
R3: Compliance with GDPR and other privacy regulations.
R4: Use of licenses like Creative Commons to clarify user rights.
R5: Assessment and documentation of data impact on society and potential biases.
Metadata Standards: Dublin Core, DataCite, schema.org
Data Curation: CKAN, DSpace, Omeka
Ethical Frameworks: Responsible AI, OpenAI Ethics Guidelines, AI4ALL
Licensing Tools: Creative Commons
Data Provenance tools: Provenance Tracking PROV-DM
3 Data Management Challenges in Large Language Models
The development of LLMs poses a complex array of data management challenges. These models engage with extensive and intricate datasets, that necessitates advanced strategies for efficient data organization and accessibility Zhao et al. (2023b). Ensuring high-quality data is crucial; this requires representative, well-curated datasets to construct models that are ethically robust and broadly applicable Bender et al. (2021). Addressing privacy and ethical concerns mandates stringent data governance to comply with ethical norms and safeguard individual privacy Wang et al. (2023). The accuracy of data annotation and labeling is also critical for model reliability Raza and Schwartz (2023), for standardized and transparent practices Monarch (2021). Additionally, the balance between data accessibility and the protection of proprietary information is key, which aligns with legal and ethical guidelines Jobin et al. (2019). Finally, adhering to data protection laws is vital for legal compliance and maintaining the ethical integrity of AI technologies, essential for preserving public confidence Raji et al. (2021). We offer a structured overview of data management challenges and requirements in LLM development in Figure 2 and summarize them below.
Figure 2: Data Management Challenges in Large Language Models
Understanding the Context: Enhanced NLP algorithms are necessary to bridge the current gap in contextual awareness Xi et al. (2023). The capacity for models to capture and interpret nuanced context is essential for advancing the intelligence and applicability of LLMs.
Accuracy and Reliability: The occasional delivery of factually incorrect responses or hallucinations Ji et al. (2023) in LLMs highlights the need for improved fact-checking mechanisms. Ensuring the accuracy and reliability of responses is crucial for the trustworthiness of these models.
Ethical and Fair Use: To combat potential biases in responses, there is a pressing need for advanced bias detection algorithms Deshpande and Sharp (2022); Wang et al. (2023); Raza et al. (2022). Ensuring the ethical application of LLMs requires a steadfast commitment to fairness and equity in model responses.
Interactivity and Personalization: Current LLMs sometimes provide generic and non-personalized responses. To address this, there is a need for adaptive learning algorithms that can learn from user interactions and provide personalized experiences He et al. (2023).
Language and Cultural Sensitivity The limited support for low-resource languages indicates a need for expanded language models and cultural datasets Chang et al. (2023). This will promote inclusivity and cultural sensitivity in LLMs.
Efficiency and Scalability: A significant gap in LLM development is the high computational resource requirement Zhao et al. (2023b). There is a need for optimization strategies that enable efficient processing without compromising the model’s performance.
Research and Development Directions Collaboration with academic institutions, ongoing model training and updates, and user-centric design approaches are essential for the continuous improvement of LLMs. This ensure that the models serve the diverse and evolving needs of users.
Mapping FAIR principles to Data Management Challenges in LLMs
: The challenges of managing data for LLMs are multifaceted, but the FAIR data principles can offer some targeted solutions. For example, Findability addresses the complexities of vast datasets by enhancing metadata and providing persistent identifiers. Interoperability standardizes data formats, enabling reliable annotation and labeling. It also plays a role in contextual awareness, improving the capacity of NLP algorithms to understand nuanced contexts. Accessibility further governs the delicate balance between sharing data and protecting proprietary information, while reusability confirms that data practices are compliant with legal and regulatory standards, ensuring that data remains usable and trustworthy. We summarize these suggestions to mapping data management challenges in LLMs to FAIR data principles in Appendix .4.
4 Framework for FAIR Data Principles Integration in LLM Development
4.1 Rationale for Integrating FAIR Data Principles in LLM Development
Our research focuses on improving Large Language Models (LLMs) by incorporating FAIR data principles into their training methods. We aim to create training datasets that adhere to FAIR principles specifically designed for LLMs. While FAIR principles have been a standard practice in AI Liang et al. (2021); Navigli et al. (2023) for a while, their integration into LLMs is a novel contribution from our team. To achieve this goal, we’ve developed a comprehensive framework that integrates FAIR principles across the entire lifecycle of LLMs, as illustrated in Figure 3. Next, we detail each phase of the LLM lifecycle, linking it specifically to our case study focused on the development of a FAIR-compliant dataset.
Figure 3: FAIR principles integrated into the LLM lifecycle
4.2 Case Study: Developing a Bias-Identification Dataset
To address the risk of LLMs perpetuating societal biases, our case study focuses on developing a dataset with a conscientious design that proactively identifies biases within the data prior to the training of these models ( whether they be LLMs or other ML models).
In this case study, “bias" is identified as a tendency within the dataset leading to the unfair representation of specific groups or outcomes Ntoutsi et al. (2020). This bias may take the form of linguistic bias Raza et al. (2024), where the dataset could unfairly favor or disfavor certain protected groups, identified by race, gender, or ethnicity, in terms of languages or dialects. It might also manifest as stereotypes Nadeem et al. (2021), toxicity Barikeri et al. (2021), or misinformation Raza and Ding (2022). Such biases typically originate from the methods used in data collection, processing, or usage. As a result, they can cause these LLMs to replicate or magnify such biases May et al. (2019). Incorporating FAIR data principles into the creation of LLMs is a proactive measure to identify and mitigate these biases before they are amplified in model outcomes. The diversity of bias dimensions and aspects addressed in our dataset is outlined in Figure 4, which shows a broad spectrum of biases including ageism, occupational jargon barriers, political rhetoric, and more.
The primary objective of this case study, aside from bias identification, is to establish a model dataset that adheres to FAIR principles. This dataset can serve as a template or example for creating datasets with various objectives for training LLMs while ensuring compliance with FAIR principles.
Figure 4: Biases across multiple dimensions explored in this study
4.2.1 Data Collection and Curation for LLM Development
FAIR Principles Achieved
: Findability, Accessibility, Interoperability, and Reusability.
In accordance with FAIR data principles, our dataset has been meticulously structured to enhance its findability through the inclusion of comprehensive metadata. This collection, sourced from diverse channels (news feeds, website) between January and May 2023, comprises over 50,000 diverse entries, significantly enhancing its accessibility and discoverability. In our efforts to engage with current social discourse, we utilized a variety of feeds and hashtags. The selected hashtags for data curation include #MediaBias, #SocialJustice, #GenderEquality, #RacialInjustice, #CulturalDiversity, #AgeismAwareness, #ReligiousTolerance, and #EconomicDisparity.
The dataset’s metadata includes information such as the dataset title, description, authors, date of creation, version, and keywords that reflect the dataset’s content and purpose. Additionally, the metadata includes keywords (LLMs, Training, Biases, News Media, NLP) for easy retrieval in these portals and specifies the data type, be it textual, numerical, or otherwise.
The Gunning Fog Index is a readability test designed to assess the level of education required to understand a piece of writing Jacobs (2023). A higher Gunning Fog Index value indicates that the text is more complex and requires a higher level of education to understand it. Figure 5 displays the distribution of Gunning Fog Index scores for our dataset, with a normal distribution centered around a mean score of 7.79. The range of scores suggests that most texts are suitable for readers with an eighth-grade education level. This data helps us evaluate the readability of these texts to match the intended audience’s comprehension abilities.
Figure 5: Histogram of the Gunning Fog Index scores for our dataset. The x-axis denotes the Gunning Fog Index scores, reflecting text complexity, and the y-axis represents the number of samples with each score.
The annotation process of this dataset is a sophisticated blend of human expertise and algorithmic efficiency, involving five domain experts and two ML-trained students. Such an annotation approach is also supported in the recent literature Liu et al. (2023); Shu et al. (2023), as being better approach compared to traditional crowd-sourcing. We employ an active learning algorithm, specifically ‘Uncertainty Sampling’ Nguyen et al. (2022) for bulk annotations. In this approach, the ML model identifies data points where human expertise can impactfully clarify ambiguities. Periodic expert reviews and inter-annotator agreement (IAA) checks, evidenced by a Cohen’s Kappa score above 0.75, ensure high-quality annotations of our data. Few IIA agreement across main bias dimensions is shown in Appendix .3.
For interoperability, the data is formatted for seamless integration for various ML tasks, such as binary and multi-label classifiers, and question answering systems, as detailed in Table 3. This enables researchers to use this data in diverse analytical contexts, facilitating cross-domain research and development. The dataset schema and examples on dataset formats are given in Appendix .2.
Table 3: Specialized Data Formats for Interoperability
Data Format Model Type
Classification Format Binary Classifier
CoNLL Format Multi-label Token Classification
SQuAD Format Question Answering System
Counterfactual Formatting Debiasing Model
Sentiment Analysis Format Sentiment Classifier
Toxicity Classification Format Toxicity Classifier
The robustness of our dataset for LLM training is established through a dual-stage quality assessment: automated analysis for initial screening of 10,000 statistically significant entries, and expert review of 500 stratified random samples for deeper insights. This method ensures the dataset’s accuracy, completeness, and consistency, crucial for LLM training, as detailed in Table 4.
Table 4: Data Quality Analysis Summary
Aspect Automated Analysis Expert Review
Accuracy 96.5% Contextually accurate
Completeness 93.0% High with occasional updates
Consistency 98.0% Uniform across samples
Furthermore, we have stored our dataset in repositories such as Huggingface, Zenodo, and Figshare, which not only adhere to metadata standards but also ensure its long-term accessibility. Overall, our data collection and curation strategy aligns with the FAIR principles and enhance the practical utility of our dataset for LLM development.
4.2.2 Model Training and Algorithm Development
FAIR Principles Achieved
: Interoperability and Reusability.
The model training and algorithm development phase is crucial for adhering to FAIR principles, particularly interoperability and reusability. We strategically deploy various LLMs, fine-tuned for tasks ranging from sentiment analysis and question answering to debiasing, and develop them with a modular design for enhanced reusability across projects.
Interoperability is achieved by using common frameworks like TensorFlow and PyTorch and standardizing data formats for inputs and outputs. Every task performed by LLMs, such as classification and debiasing, is executed following the preparation of specific training formats tailored to each task. This ensures seamless integration of our models with various systems and datasets. We further enhance this with the inclusion of model cards for each LLM, providing essential information such as model purpose, performance, and usage guidelines, thus aiding in understanding and adoption.
We prepare API documentation that provides details for every aspect of training and development, including parameter settings and algorithm modifications. This transparency facilitates scientific validation and progress Jacobsen et al. (2020b). We also foster community collaboration by sharing our findings, models, and tools. The goal is to contribute to public repositories, and participating in AI research platforms.
4.2.3 Model Evaluation and Validation
FAIR Principles Achieved
: Reusability and Accessibility.
In our evaluation and validation phase, we prioritize the FAIR principles of Reusability and Accessibility to ensure that our models and findings can be widely utilized and understood. Transparency in reporting and the provision of accessible documentation are key aspects of our methodology. All results, including bias analysis and model benchmarking, are made publicly available, allowing for comprehensive observation and application by the wider research community. An example of this approach is presented in our detailed bias analysis heatmap (Figure 6) and the benchmarking table (Table 5), which illustrate our commitment to these principles.
Figure 6: Heatmap of labels across main dimensions
In Table 5, we present performance metrics across various tasks, demonstrating the effectiveness of our classifiers in toxicity detection, bias classification, sentiment analysis, multi-label token classification, and QA capabilities. Additionally, we integrated a debiasing process using LLaMa 2B Chat Touvron et al. (2023), re-evaluating our models on the same test set post-debiasing. This allowed us to assess the impact of debiasing on reducing toxicity and bias, and its influence on sentiment analysis. The debiasing process notably improved scores and indicated the efficacy of the debiasing intervention.
Table 5: Comprehensive Analysis of Language Model Performance and Impact of Debiasing. The first section details performance metrics of various models on different tasks, fine-tuned for optimal settings. The second section highlights the effect of debiasing on toxicity, bias classification, and sentiment scores.
Performance Metrics of Language Models. Higher the score, better the performance.
Model (Type) Accuracy Precision F1-Score
Binary Bias Classifier (BERT-large-uncased) 92% 89% 90%
NER Model (RoBERTa-large-uncased) 95% 93% 94%
QA System (BERT-base-uncased) 88% 87% 86%
Sentiment Classifier (BERT-large-uncased) 90% 91% 90%
Toxicity Classifier (BERT-large-uncased) 91% 90% 91%
LLaMa2 (7b-chat-hf) 93% 94% 93%
Impact of Debiasing. Lower scores post-debiasing is better.
Metric Pre-Debiasing Score Post-Debiasing Score
Toxicity Score 75% 24%
Bias Classification Score 80% 55%
Sentiment Score -0.3 0.2
The Table 5 demonstrates that all models initially exhibit strong classification performance. However, after the debiasing process, a noticeable decline in classification performances is observed across these models when applied to the debiased dataset. This suggests that the original training was heavily reliant on biased data, and the introduction of debiased data disrupts the model’s classification ability, leading to a reduction in accuracy. This result aligns very much with the recent safety benchmarking in recent works Touvron et al. (2023), which shows there is trade-off highlights the impact of debiasing on model performance. Overall,these models’ evaluation and validation phase, emphasize transparent reporting and public availability of models and the detailed results, aligns with the FAIR principles of Reusability and Accessibility.
4.2.4 Deployment and Ongoing Monitoring
FAIR Principles Achieved
: Accessibility, Reusability, and Findability.
In the post-deployment phase, we emphasize the FAIR principles of Accessibility, Reusability, and Findability. Our approach includes providing comprehensive documentation and clear usage guidelines to enhance the model’s findability and accessibility. We employ version control systems to ensure easy tracking and retrieval of the latest model versions, supporting both findability and reusability.
To facilitate user engagement and model effectiveness, we offer extensive training materials and support, making the models more accessible to a diverse user base. Regular reporting of performance metrics offers insights into the models’ ongoing effectiveness and areas for improvement, aiding in their reusability. We also provide integration support with various tools and platforms, ensuring that our models can be easily adopted in different technological contexts.
These practices collectively ensure that our deployed models adhere to the FAIR principles, maintaining their utility and efficacy in a dynamic, user-centric environment.
4.2.5 Community Engagement and Collaborative Development
FAIR Principles Achieved
: Findability, Accessibility, Interoperability, and Reusability.
In our efforts to foster community engagement and collaborative development, we align with the FAIR principles of Findability, Accessibility, Interoperability, and Reusability. We provide open-source development to facilitate collaborative innovation. Our datasets are provided in accessible formats such as JSON and CSV and are indexed with rich metadata. This approach ensures that our LLMs are developed in environments conducive to collaborative progress.
The datasets are disseminated under the Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license Creative Commons (2023), aligning our work with the FAIR principles of data stewardship.
Upholding FAIR Principles
We provide a comprehensive analysis in Table 6, showcasing our commitment to upholding the FAIR data principles to enhance the dataset’s usefulness for both present and future research endeavors.
Table 6: FAIR Compliance in Bias Identification Dataset and LLM Development
FAIR Dimension FAIR Compliance Criteria Met Remarks (Implementation Details)
Findability
F1: Rich, descriptive metadata
F2: Standardized data indexing
F3: Documented data sources
F4: Persistent identifiers for data and models
F1: Metadata includes descriptions for data and models.
F2: Data and models indexed with relevant keywords.
F3: Clear documentation of data sources and model development.
F4: Unique identifiers for data and models.
Accessibility
A1: Clearly defined access protocols
A2: Long-term data and model preservation
A3: Data and models available in accessible formats
A1: Open-access repository for data and models.
A2: Long-term preservation of data and models.
A3: Data in CSV, JSON, XML; Models accessible via APIs.
Interoperability
I1: Use of standard data and model formats
I2: Adherence to data and model exchange standards
I1: Data in standard formats; Models compatible with common ML frameworks.
I2: Adheres to standard data and model exchange protocols.
Reusability
R1: Detailed metadata for data and models
R2: Adherence to ethical and transparency standards
R3: Licensing information for data and model usage
R1: Metadata includes usage guidelines for data and models.
R2: Focus on ethical data use and transparent model training.
R3: Data licensed under $C$ BY-NC 4.0; Models with open-source licenses.
This approach ensures that our LLMs are technologically advanced, ethically sound, and conducive to a robust, inclusive future in AI development.
5 Discussion
5.1 Limitations of the FAIR Data Principles in Addressing Data Challenges
The FAIR data principles are a crucial step towards more open and efficient use of data in research and beyond. However, their effectiveness is limited by challenges related certain factors discussed below:
  • Data Quality Oversight: Although the FAIR principles make data more accessible and usable, these princicples do not inherently guarantee the quality or validity of the data. There is no built-in mechanism to ensure the data’s accuracy, completeness, or reliability. This lack of quality control can lead to the proliferation of poor-quality data and can impact research and decision-making processes.
  • Ethical and Privacy Concerns: While FAIR principles promote accessibility and sharing, they do not sufficiently address the complex ethical implications and privacy issues related to sensitive data. In areas such as patient health records or personal information, the unrestricted sharing of data could lead to privacy violations and ethical dilemmas.
  • Resource and Cost Implications: Implementing FAIR principles can be resource-intensive. It requires infrastructure, tools, and expertise to make the process happen. This can be particularly challenging for smaller institutions or developing countries with limited resources, potentially widening the digital divide.
  • One-Size-Fits-All Approach: The FAIR principles is not one-size-fits-all solution for all scientific disciplines. Different fields may require tailored approaches to data management that the generalized FAIR guidelines do not fully accommodate.
  • Lack of Universally Accepted Standards: While the FAIR principles provide a framework, there is still a lack of universally accepted standards and metrics to evaluate how ’FAIR’ a dataset is. This ambiguity can lead to inconsistent applications of the principles.
  • Technical Challenges in Interoperability: Data from different sources often exist in various formats and standards, it is difficult to make seamless integration across different systems and platforms. The over-emphasis on data sharing might overshadow other important aspects of data stewardship, such as data privacy.
  • Intellectual Property and Commercial Interests: FAIR principles may conflict with intellectual property rights and commercial interests. Making data freely accessible and reusable might undermine proprietary research or commercial advantages, leading to resistance from certain sectors.Adhering to the FAIR principles must also be balanced with compliance with legal and regulatory frameworks, which can vary significantly across regions.
  • Training and Awareness: Effective implementation of FAIR principles requires a high level of awareness and training among data practitioners. The lack of such training can be a significant barrier to the adoption of these principles.
5.2 Strategies for Mitigating Limitations and Enhancing Data Utility
In addressing the limitations of FAIR Data Principles and enhancing data utility, a multifaceted strategy is essential. This involves refining data quality through rigorous validation processes, ensuring accuracy and reliability. Simultaneously, balancing data accessibility with ethical and privacy considerations is crucial, there is need to employ advanced anonymization techniques and strict privacy protocols. Emphasizing data preservation and sustainability alongside sharing, broadens the scope of stewardship. Integrating legal and regulatory compliance is vital for the robust and ethical utilization of data in research and practice.
5.3 Limitation of the Study and Future Perspectives
Our Bias-Identification dataset, designed to identify and mitigate known biases, may inadvertently overlook emerging biases, thereby emphasizing the need for ongoing revision and vigilant monitoring May et al. (2019); Raza (2021). In tandem with the rapid development of LLMs, this necessitates adaptable datasets and methodologies to accommodate new architectures and ensure wide-ranging applicability Wilson et al. (2023). Concurrently, scaling the dataset to match the complexity of advanced LLMs, while adhering to ethical standards, emerges as a substantial challenge. This endeavor is compounded by the imperative to mitigate biases in model interpretations, bolster interpretability, and ensure robust data privacy and security Zhao et al. (2023a); Chen et al. (2023). Therefore, future research directions should consider creating dynamic mechanisms for dataset updates, advancing bias mitigation techniques, aligning datasets with evolving LLM technologies, exploring scalable dataset maintenance solutions, and broadening the spectrum of LLM applications. Integral to this journey is the development of comprehensive ethical guidelines and the fostering of collaborative research, both of which are pivotal for the responsible evolution of LLM technology.
6 Conclusion
Our study integrates FAIR data principles into LLM development, enhancing data management and model training. This has led to a robust and versatile dataset, useful in various research contexts and offering insights into applying FAIR principles in LLM training. Despite advancements, there are limitations and opportunities for future research. These include updating datasets to reflect emerging trends, improving bias detection, adapting to new LLM architectures, enhancing model interpretability, and developing ethical guidelines. Our work contributes to the responsible development of AI, aiming to create more ethical and effective AI tools that benefit diverse communities.
Research Ethics and Social Impact
Ethical Considerations Statement
In conducting this research, we carefully adhered to ethical guidelines to ensure the integrity and societal value of our work. We ensured the privacy and confidentiality of data, obtaining consent where required, and employed measures to safeguard sensitive information. Our methodology was designed to minimize any potential harm, with a focus on responsible data handling and analysis. Recognizing the potential implications of our research, we took steps to ensure that our findings are presented in a manner that is respectful and considerate of diverse perspectives and experiences.
Researcher Positionality Statement
Our team’s diverse academic and cultural backgrounds significantly influenced our approach to this research. Our team comprises members from both academia and industry, bringing a wealth of different viewpoints to the study. This diversity enriched our analysis, allowing for a more comprehensive understanding of the data and its implications. We believe that acknowledging our positionality is crucial for maintaining transparency and objectivity in our research.
Adverse Impact Statement
We recognize the potential adverse impacts of our research, particularly given the sensitive nature of the dataset prepared for bias identification. While our aim is to enhance fairness and inclusivity, there is a risk that the data or findings could be misinterpreted or misapplied in a manner that inadvertently negatively portrays underrepresented groups. We are committed to ensuring our case study does not contribute to further stigmatization or discrimination of any group. We emphasize the importance of using our results responsibly, particularly in contexts that were not directly intended in our study, to avoid potential ethical dilemmas or societal harm.
Appendices
.1 Acronyms and Full Forms
Academia.edu, Academia.edu; AI4ALL, AI4ALL; Algolia, Algolia; Amazon S3, Amazon Simple Storage Service; Apache Atlas, Apache Atlas; Apache Lucene, Apache Lucene; Apache NiFi, Apache NiFi; Apache Solr, Apache Solr; Archivematica, Archivematica; CKAN, Comprehensive Knowledge Archive Network; Clockss, Controlled LOCKSS; Collibra, Collibra; Creative Commons, Creative Commons; Crossref, Crossref; DSpace, DSpace; Dublin Core, Dublin Core Metadata Initiative; Data Provenance Tools, Data Provenance Tools; DataCite, DataCite; Dataverse, Microsoft Dataverse; EPrints, EPrints; Ecoinformatics, Ecoinformatics; Elasticsearch, Elasticsearch; FGED, Functional Genomics Data Society; Figshare, Figshare; GDPR, General Data Protection Regulation; Google Cloud Storage, Google Cloud Storage; GraphQL, GraphQL; HL7 FHIR, Health Level Seven Fast Healthcare Interoperability Resources; IEDB, Immune Epitope Database; ISO/IEC, International Organization for Standardization/International Electrotechnical Commission; LOCKSS, Lots of Copies Keep Stuff Safe; NCBI, National Center for Biotechnology Information; OAI-PMH, Open Archives Initiative Protocol for Metadata Harvesting; Omeka, Omeka; OneTrust, OneTrust; OpenAI Ethics Guidelines, OpenAI Ethics Guidelines; OpenAPI, OpenAPI Initiative; OpenRefine, OpenRefine; ORCID, Open Researcher and Contributor ID; OWL, Web Ontology Language; Portico, Portico; PROV-DM, Provenance Data Model; RDF, Resource Description Framework; RE3data, Registry of Research Data Repositories; ResearchGate, ResearchGate; Responsible AI, Responsible AI; REST, Representational State Transfer; schema.org, Schema.org; SOAP, Simple Object Access Protocol; SPARQL, SPARQL Protocol and RDF Query Language; Talend, Talend; TrustArc, TrustArc; XOD, eXtensible ontology development; XSLT, Extensible Stylesheet Language Transformations; Zenodo, Zenodo; gRPC, gRPC Remote Procedure Calls.
.2 Dataset Schema and Formats
The dataset has multiple attributes like Text, Dimension, Biased Words, Aspect, Bias Label, Sentiment, Toxic, Identity Mention, Debasied Text.
"Text": "Lawyers are always manipulative and cannot be trusted.",
"Dimension": "Professional Integrity",
"Biased Words": "always, manipulative, cannot be trusted",
"Aspect": "Trustworthiness of lawyers",
"Bias Label": "BIASED",
"Sentiment": "Negative",
"Toxic": "Yes",
"Identity Mention": "Lawyers",
"Debiased Text": "Trustworthiness is an individual trait and varies among professionals, including lawyers."
We provide examples for each data format, demonstrating the use of dataset attributes in diverse machine learning models:
Table 1: Examples of Dataset Formats
Format
Example
Classification
{ "Text": "Politicians are often corrupt.", "Bias Label": "BIASED", "Sentiment": "Negative" }
CoNLL
{ "Text": "Doctors are always caring.", "Biased Words": ["always", "caring"], "Identity Mention": "Doctors" }
SQuAD
{ "Text": "Why are lawyers untrustworthy?", "Aspect": "Trustworthiness", "Debiased Text": "Trustworthiness varies individually." }
Counterfactual
{ "Text": "Young people are irresponsible with money.", "Biased Words": ["irresponsible"], "Debiased Text": "Financial habits vary by individual." }
Sentiment/Toxicity
{ "Text": "Politicians are dishonest.", "Sentiment": "Negative", "Toxic": "Yes" }
.3 Inter Annotators Agreement
Figure1 illustrates the expert consensus aligned with active learning outcomes across various bias dimensions, showing the highest agreement in ’Socioeconomic Bias’.
Figure 1: Expert Agreement Across Bias Dimensions. The bar graph quantifies the concordance between domain experts evaluations and the model’s predictions.
.4 Data Management Challenges in LLMs and Corresponding FAIR Principles
Table 2: Data Management Challenges in LLMs and Corresponding FAIR Principles
Data Management Challenge FAIR Principle Addressed
Vast and Complex Datasets
Findability through detailed metadata and persistent identifiers to enhance data discovery .
Data Quality and Bias
Accessibility with ethical access protocols to provide high-quality, unbiased data.
Privacy and Ethical Concerns
Reusability with clear legal and ethical documentation to uphold privacy and ethics.
Data Annotation and Labeling
Interoperability through standardized data formats for reliable annotation and labeling.
Data Accessibility and Sharing
Accessibility to ensure a balance between open data sharing and protection of proprietary information .
Legal and Regulatory Compliance
Reusability to align data management with legal standards for future use .
Contextual Awareness
Interoperability for enhanced NLP algorithms to accurately capture and interpret context.
Accuracy and Reliability
Accessibility and Reusability to ensure mechanisms for improved fact-checking and consistent data quality.
Ethical and Fair Use
Accessibility with advanced bias detection algorithms to promote fairness .
Interactivity and Personalization
Findability and Accessibility for adaptive learning from user interactions for personalized experiences.
Language and Cultural Sensitivity
Interoperability to support expanded language models and cultural datasets for inclusivity .
Technical and Scalability
Reusability for efficient processing strategies that facilitate the scalable use of computational resources.